A photon of wavelength 300 mm is absorbed by a gas and then re-emitted as two photos, one is red with wavelength 760 nm. The wave number of second photon will be?

Grade:11

1 Answers

Vikas TU
7 years ago
Conserve the monetum here.
k = h/p = h /mv.......................................(1)
Therefore,
M*v = m1v1 + m2v2
or
h/k = h/k1 + h/k2 (from eqn1. )
or 
1/k = = 1/k1 + 1/k2
now substitute the main wavelengths to get the values Hence.